# Cartesian and spherical multipole expansions in anisotropic media

### v2

@@ Abstract
The multipole expansion can be formulated in spherical and Cartesian coordinates. By constructing an explicit map linking both formulations in isotropic media, we discover a lack of equivalence between them in anisotropic media. In isotropic media, the Cartesian multipole tensor can be reduced to a spherical tensor containing fewer entries. In anisotropic media, however, the loss of propagation symmetry prevents this reduction. Consequently, non-harmonic sources radiate fields that can be projected onto a finite set of Cartesian multipole moments but require potentially infinitely many spherical moments. For harmonic sources, the link between the two approaches provides a systematic way to construct the spherical multipole expansion from the Cartesian one. The lack of equivalence between both approaches results in physically significant effects wherever the field propagation includes the Laplace operator. We demonstrate this issue in an electromagnetic radiation inverse problem in anisotropic media, including an analysis of a large-anisotropy regime and an introduction to vector spherical harmonics. We show that the use of the Cartesian approach increases the efficiency and interpretability of the model. The proposed approach opens the door to wider applications of the multipole expansion and new degrees of freedom in anisotropic media.
@@ I Introduction
The multipole expansion method has long been an essential mathematical tool successfully applied to a wide range of systems described by partial differential equations involving the Laplace operator. Commonly used in electromagnetism [1, 2], this method also sees applications from the quantum level [3] up to Newtonian gravitational interaction [4, 5] and gravitational waves [6, 7]; indeed, whenever one needs to describe how fields behave outside the source region. Notably, the multipole expansion in anisotropic media, where wave propagation becomes direction-dependent, remains relatively ill-studied. This must be distinguished from the multipole expansion associated with an anisotropic source, which results in direction-dependent radiation or scattering [8]. Here, we assume that wave propagation occurs within an anisotropic medium, while the source itself may also exhibit anisotropic characteristics.
For optical or radio-frequency systems, the rise of advanced wave manipulation, such as via metamaterials [9, 10, 11] or nanophotonics [12], warrants a study of this method in anisotropic media. For instance, meta-atoms mixing electric and magnetic multipole moments give rise to bianisotropy [13]. The multipole expansion is also readily applied to metasurfaces to express the fields, which are then linked through the $T$-matrix method [14]. This has been applied, for example, to the analysis of substrate-supported dielectric nanoresonator metasurfaces [15]. The multipole expansion is also used to obtain a basis of resonant states in the resonant-state expansion method for the analysis of open optical systems [16]. High-order multipoles are also gaining interest in connection to the chosen center of coordinates [17]. A variation of the multipole expansion can also be used to describe the nonlocal response of certain metasurfaces [18]. While the literature commonly uses the multipole expansion to describe and analyze metamaterials with anisotropic wave propagation, the analysis of the multipole expansion of fields embedded in an anisotropic medium is missing. Implications of this analysis could include the effect of a layered substrate with uniaxial wave propagation on the interaction between neighboring particles in a metasurface.
@@
The rest of the paper is structured as follows. In Section II, we describe the link between the spherical and Cartesian multipole expansions in isotropic media and demonstrate this equivalence breaks down in anisotropic media. Next, we also discuss the link to the vector spherical harmonics: the extension of their scalar counterpart to vector fields; the standard for considering direction-dependent wave propagation. In Section III, we illustrate the lack of equivalence in a radiation inverse problem, including a detailed analysis of a large-anisotropy regime. Finally, we conclude with some remarks.
II Multipole expansions in anisotropic media
@@ II.1 Spherical and Cartesian multipole expansions in isotropic media
We begin by finding equivalence relations between the spherical and Cartesian multipole expansions by analyzing their singular behavior. This will allow us to extrapolate the approach to anisotropic media by considering the well-defined radiation of point sources. To this end, previous work on the pseudopotential method [28, 29, 30] has sparked interest in the singular behavior of spherical wave functions. This behavior is challenging to describe as the concept of “radial” derivatives of the Dirac $\delta$ function is not trivially defined (see [31, 32, 33]) because of the singular nature of spherical coordinates at the origin. A rigorous approach in [34] allows to write, for the spherical basis function $f_{\ell m}(\mathbf{r})$ introduced below,
$(\nabla^{2}+k_{0}^{2})f_{\ell m}(\mathbf{r})=\frac{(-1)^{\ell}4\pi i}{k_{0}^{\ell+1}}Y_{\ell m}(\nabla)\delta(\mathbf{r})=:\tilde{Y}_{\ell m}(\nabla)\delta(\mathbf{r})$ (1)
@@
$\mathcal{Y}_{\ell m}(\theta,\phi)=\sqrt{\frac{2\ell+1}{4\pi}}\sqrt{\frac{(l-m)!}{(l+m)!}}P_{\ell m}\left[\cos(\theta)\right]e^{im\phi}$ (2)
@@
and likewise for $\tilde{Y}_{\ell m}(\nabla)$. For instance, $Y_{1,1}(\mathbf{k})=-1/2\sqrt{3/(2\pi)}(k_{x}+ik_{y})$, hence $Y_{1,1}(\nabla)\delta(\mathbf{r})=-1/2\sqrt{3/(2\pi)}(\partial\delta(\mathbf{r})/\partial x+i\partial\delta(\mathbf{r})/\partial y)$.
@@
vp denotes the principal value. This generalizes to higher orders as [34, Lemma 1.(c)]
@@
Radiating waves are described by spherical Hankel functions of the first or second kind $h_{\ell}^{(1,2)}(k_{0}r)$ [35, Eq. 10.1.1]:
$h_{\ell}^{(1)}(k_{0}r)=j_{\ell}(k_{0}r)+iy_{\ell}(k_{0}r)$ (7)
@@
$f_{\ell m}(\mathbf{r})=\text{vp}\left[h_{\ell}^{(1)}(k_{0}r)\mathcal{Y}_{\ell m}(\theta,\phi)\right]\iff\hat{Y}(\mathbf{k})$ (13)
@@ II.1.1 Spherical to Cartesian multipole expansions
$\tilde{Y}_{2,0}(\mathbf{k})=i\sqrt{5\pi}/k_{0}^{3}(2k_{z}^{2}-k_{x}^{2}-k_{y}^{2})$ (15)
@@
$f_{2,0}(\mathbf{r})=i\sqrt{5\pi}/k_{0}^{3}\left[2f_{(0,0,2)}(\mathbf{r})-f_{(2,0,0)}(\mathbf{r})-f_{(0,2,0)}(\mathbf{r})\right].$ (16)
@@ II.1.2 Cartesian to spherical multipole expansions
Surprisingly, it is generally not possible to write an arbitrary monomial $\mathbf{k}^{\alpha}$ as a linear combination of harmonic polynomials $\tilde{Y}_{\ell m}(\mathbf{k})$, as the monomial is not necessarily harmonic. For any monomial, we can nonetheless write by [38, Theorem 5.21]:
@@
$\tilde{A}_{|\alpha|-2n,m}=\frac{k_{0}^{\ell+1}}{(-1)^{\ell}4\pi i}A_{|\alpha|-2n,m}$ (23)
@@
$k_{z}^{2}=-ik_{0}/6\left[1/\sqrt{\pi}r^{2}\tilde{Y}_{0,0}(\mathbf{k})+2k_{0}^{2}/\sqrt{5\pi}\tilde{Y}_{2,0}(\mathbf{k})\right]$ (29)
@@
$f_{(0,0,2)}(\mathbf{r})=ik_{0}^{3}/6\left[1/\sqrt{\pi}f_{0,0}(\mathbf{r})-2/\sqrt{5\pi}f_{2,0}(\mathbf{r})\right]$ (30)
@@ II.2 The case of anisotropic media
In anisotropic media, the Laplace operator (as in Equation (25)) is replaced with weighted second-order derivatives, reflecting different propagation characteristics along different directions. For example, electromagnetic fields obey a modified vector wave equation [39, Equation (5.6)] whose anisotropy is determined by the constitutive parameters’ tensorial nature. In such cases, as we illustrate below, some sources radiate fields that cannot be fully represented by a spherical multipole expansion. In turn, the more numerous Cartesian moments are necessary to model these fields accurately and efficiently.
@@ II.3 Relation to the vector spherical harmonics
Many applications such as optics require the use of vector spherical harmonics. Here, we wish to outline how the link between the spherical and Cartesian approaches can be extended to the vector spherical harmonics and, thus, the vector wave equation. Fortunately, it is possible to write the components of vector spherical harmonics in terms of scalar spherical harmonics. Using quantum theory formalism and the Condon–Shortley phase, the vector spherical harmonics can be written as [40, Eq. 7.3.3]
@@
$\mathbf{e}_{\pm 1}=\mp\frac{1}{\sqrt{2}}\left(\mathbf{e}_{x}\pm i\mathbf{e}_{y}\right)$ (32)
@@
The vector spherical harmonics used by Jackson [24] $\bm{\mathcal{Y}}_{\ell m}^{(q)}(\theta,\phi)$, $q=0,\pm 1$, are obtained as linear combinations of $\bm{\mathcal{Y}}^{\ell}_{jm}(\theta,\phi)$ (see Section I of the Supplementary Information). They correspond to the azimuthal ($q=0$), polar ($q=1$), and radial ($q=-1$) components. In homogeneous media, the radial component is not allowed for electromagnetic fields. Together with Equation (31), we show in the same section of the Supplementary Information how to express the azimuthal or polar vector spherical harmonics using the Cartesian multipole expansion of an appropriate current source. We show that a current density given by
@@
where $\mathbf{Y}_{\ell\pm 1,m}^{\ell}(\mathbf{k})=k^{\ell}\bm{\mathcal{Y}}_{\ell\pm 1,m}^{\ell}(\mathbf{k})$. The Cartesian components of the current densities $\mathbf{J}_{\ell m}^{(0,1)}(\mathbf{r})$ correspond to singularities of the scalar wave equation analyzed above. Hence, any vector spherical harmonic can be represented by an appropriate sum of Cartesian multipoles.
As expected from the findings for the scalar wave equation, the inverse transformation—i.e., writing an arbitrary Cartesian multipole expansion as a linear combination of vector spherical harmonics—is not as straightforward, and a complete treatment goes beyond the scope of this research. To understand the issue at hand, it is enlightening to count the number of vector spherical harmonics at a given order. The azimuthal vector harmonics correspond to a current density given by the vector spherical harmonics $\bm{\mathcal{Y}}_{\ell m}^{\ell}(\theta,\phi)$. In light of Equation (31), there are $12\ell+$ 1 different azimuthal waves of order $\ell>0$ (none for $\ell=0$). By the same argument, there are $12(\ell\pm 1)+$ 1 polar vector harmonics of order $\ell>0$ corresponding to $j=\ell\pm 1$. But the polar harmonics are degenerate, in that both values of $j$ yield the same waves, by virtue of the contraction in Equation (26) (see the Supplementary Information for a detailed derivation). Thus, in total, there are
@@
However, since there are three possible polarizations ($x$, $y$, and $z$) and $12\ell+$ 1 spherical harmonics of order $\ell$ per polarization, one would expect
@@
different vector spherical harmonics. The missing harmonics are the degenerate polar harmonics with $j=\ell-1$, and their count is given by 2($12(\ell-1)+$ 1:
$\underbrace{4\ell+4}_{\text{actual}}+\underbrace{2(\ell-1)+1}_{\text{degenerate}}=\underbrace{6\ell+3}_{\text{expected}}$ (38)
This hints towards a similar issue to the scalar case when going to anisotropic media, as the Laplace operator is responsible for the degeneracy. Another viewpoint is that in anisotropic media, the degeneracy disappears, and novel degrees of freedom in wave polarization become available.
To note, we circumvent this issue in the application below by opting for the electric Green’s tensor-valued function instead of vector spherical harmonics. This approach enables a direct connection between the electric current and the vector-valued field.

### v3

@@ Abstract
The multipole expansion can be formulated in spherical and Cartesian coordinates. By constructing an explicit map linking both formulations in isotropic media, we discover a lack of equivalence between them in anisotropic media. In isotropic media, the Cartesian multipole tensor can be reduced to a spherical tensor containing fewer independent components. In anisotropic media, however, the loss of propagation symmetry prevents this reduction. Consequently, non-harmonic sources radiate fields that can be projected onto a finite set of Cartesian multipole moments but require potentially infinitely many spherical moments. For harmonic sources, the link between the two approaches provides a systematic way to construct the spherical multipole expansion from the Cartesian one. The lack of equivalence between both approaches results in physically significant effects wherever the field propagation includes the Laplace operator. We demonstrate this issue in an electromagnetic radiation inverse problem in anisotropic media, including an analysis of a large-anisotropy regime and an introduction to vector spherical harmonics. We show that the use of the Cartesian approach increases the efficiency and interpretability of the model. The proposed approach opens the door to a broader application of the multipole expansion in anisotropic media.
@@ I Introduction
The multipole expansion method has long been an essential mathematical tool successfully applied to a wide range of systems described by partial differential equations involving the Laplace operator. Commonly used in electromagnetism [2, 3], this method also sees applications from the quantum level [4] up to Newtonian gravitational interaction [5, 6] and gravitational waves [7, 8]; indeed, whenever one needs to describe how fields behave outside the source region. However, the multipole expansion in anisotropic media, where wave propagation becomes direction-dependent, has remained relatively ill-studied. This must be distinguished from the multipole expansion associated with an anisotropic source, which results in direction-dependent radiation or scattering [9]. Here, we assume that wave propagation occurs within an anisotropic medium, while the source itself may also exhibit anisotropic characteristics.
For optical or radio-frequency systems, the rise of advanced wave manipulation, such as via metamaterials [10, 11, 12] or nanophotonics [13], warrants a study of this method in anisotropic media. For instance, meta-atoms mixing electric and magnetic multipole moments give rise to bianisotropy [14]. The multipole expansion is also readily applied to metasurfaces to express the fields, which are then linked through the $T$-matrix method [15]. For example, this approach has been applied to the analysis of substrate-supported dielectric nanoresonator metasurfaces [16]. The multipole expansion is also used to obtain a basis of resonant states in the resonant-state expansion method for the analysis of open optical systems [17]. High-order multipoles are also gaining interest in connection to the chosen center of coordinates [18]. A variation of the multipole expansion can also be used to describe the nonlocal response of certain metasurfaces [19]. While the literature commonly uses the multipole expansion to describe and analyze metamaterials with anisotropic wave propagation, the analysis of the multipole expansion of fields embedded in an anisotropic medium is missing. Implications of this analysis could include the effect of a layered substrate with uniaxial wave propagation on the interaction between neighboring particles in a metasurface.
@@
In this paper, we describe the link between the spherical and Cartesian multipole expansions in isotropic media and demonstrate this equivalence breaks down in anisotropic media. Next, we also discuss the link to the vector spherical harmonics: the extension of their scalar counterpart to vector fields. We then illustrate the lack of equivalence in a radiation inverse problem, including a detailed analysis of a large-anisotropy regime. We show that, in anisotropic media, the use of the Cartesian multipole expansion improves the modeling efficiency and interpretability.
II Methods
@@ II.1 Spherical and Cartesian multipole expansions in isotropic media
We begin by finding equivalence relations between the spherical and Cartesian multipole expansions by analyzing their singular behavior. This will allow us to extrapolate the approach to anisotropic media by considering the well-defined radiation of point sources. To this end, previous work on the pseudopotential method [29, 30, 31] has sparked interest in the singular behavior of spherical wave functions. This behavior is challenging to describe as the concept of “radial” derivatives of the Dirac $\delta$ function is not trivially defined (see [32, 33, 34]) because of the singular nature of spherical coordinates at the origin. A rigorous approach in [35] allows to write, for the spherical basis function $f_{\ell m}(\mathbf{r})$ introduced in Equation (8),
$(\nabla^{2}+k_{0}^{2})f_{\ell m}(\mathbf{r})=\frac{(-1)^{\ell}4\pi\text{i}}{k_{0}^{\ell+1}}Y_{\ell m}(\nabla)\delta(\mathbf{r})=:\tilde{Y}_{\ell m}(\nabla)\delta(\mathbf{r})$ (1)
@@
$\mathcal{Y}_{\ell m}(\theta,\phi)=\sqrt{\frac{2\ell+1}{4\pi}}\sqrt{\frac{(l-m)!}{(l+m)!}}P_{\ell m}\left[\cos(\theta)\right]\text{e}^{\text{i}m\phi}$ (2)
@@
and likewise for $\tilde{Y}_{\ell m}(\nabla)$. For instance, $Y_{1,1}(\mathbf{k})=-1/2\sqrt{3/(2\pi)}(k_{x}+\text{i}k_{y})$, hence $Y_{1,1}(\nabla)\delta(\mathbf{r})=-1/2\sqrt{3/(2\pi)}(\partial\delta(\mathbf{r})/\partial x+\text{i}\partial\delta(\mathbf{r})/\partial y)$.
@@
vp denotes the principal value. This generalizes to higher orders as [35, Lemma 1.(c)]
@@
Radiating waves are described by spherical Hankel functions of the first or second kind $h_{\ell}^{(1,2)}(k_{0}r)$ [36, Eq. 10.1.1]:
$h_{\ell}^{(1)}(k_{0}r)=j_{\ell}(k_{0}r)+\text{i}y_{\ell}(k_{0}r)$ (7)
@@
$f_{\ell m}(\mathbf{r})=\text{vp}\left[h_{\ell}^{(1)}(k_{0}r)\mathcal{Y}_{\ell m}(\theta,\phi)\right]\iff\tilde{Y}(\mathbf{k})$ (13)
@@ II.1.1 Spherical to Cartesian multipole expansions
$\tilde{Y}_{2,0}(\mathbf{k})=\text{i}\sqrt{5\pi}/k_{0}^{3}(2k_{z}^{2}-k_{x}^{2}-k_{y}^{2})$ (15)
@@
$f_{2,0}(\mathbf{r})=\text{i}\sqrt{5\pi}/k_{0}^{3}\left[2f_{(0,0,2)}(\mathbf{r})-f_{(2,0,0)}(\mathbf{r})-f_{(0,2,0)}(\mathbf{r})\right].$ (16)
@@ II.1.2 Cartesian to spherical multipole expansions
Nevertheless, it is generally not possible to write an arbitrary monomial $\mathbf{k}^{\alpha}$ as a linear combination of harmonic polynomials $\tilde{Y}_{\ell m}(\mathbf{k})$, as the monomial is not necessarily harmonic. For any monomial, we can nonetheless write by [39, Theorem 5.21]:
@@
$\tilde{A}_{|\alpha|-2n,m}=\frac{k_{0}^{\ell+1}}{(-1)^{\ell}4\pi\text{i}}A_{|\alpha|-2n,m}$ (23)
@@
$k_{z}^{2}=-\text{i}k_{0}/6\left[1/\sqrt{\pi}k^{2}\tilde{Y}_{0,0}(\mathbf{k})+2k_{0}^{2}/\sqrt{5\pi}\tilde{Y}_{2,0}(\mathbf{k})\right]$ (29)
@@
$f_{(0,0,2)}(\mathbf{r})=\text{i}k_{0}^{3}/6\left[1/\sqrt{\pi}f_{0,0}(\mathbf{r})-2/\sqrt{5\pi}f_{2,0}(\mathbf{r})\right]$ (30)
@@ II.2 The case of anisotropic media
In anisotropic media, the Laplace operator (as in Equation (25)) is replaced with weighted second-order derivatives, reflecting different propagation characteristics along different directions. For example, electromagnetic fields obey a modified vector wave equation [40, Equation (5.6)] whose anisotropy is determined by the constitutive parameters’ tensorial nature. In such cases, as we illustrate in the section “Illustration on an inverse problem,” some sources radiate fields that cannot be fully represented by a spherical multipole expansion. In turn, the more numerous Cartesian moments are necessary to model these fields accurately and efficiently.
@@ II.3 Relation to the vector spherical harmonics
Many applications such as optics require the use of vector spherical harmonics. Here, we wish to outline how the link between the spherical and Cartesian approaches can be extended to the vector spherical harmonics and, thus, the vector wave equation.
II.3.1 From vector spherical harmonics to the Cartesian multipole expansion
Fortunately, it is possible to write the components of vector spherical harmonics in terms of scalar spherical harmonics. This allows to generalize the reasoning applied to the scalar wave equation. Using quantum theory formalism and the Condon–Shortley phase, the vector spherical harmonics can be written as [41, Eq. 7.3.3]
@@ II.3.1 From vector spherical harmonics to the Cartesian multipole expansion
$\mathbf{e}_{\pm 1}=\mp\frac{1}{\sqrt{2}}\left(\mathbf{e}_{x}\pm\text{i}\mathbf{e}_{y}\right)$ (32)
@@
The vector spherical harmonics used by Jackson [25] $\bm{\mathcal{Y}}_{\ell m}^{(q)}(\theta,\phi)$, $q=0,\pm 1$, are obtained as linear combinations of $\bm{\mathcal{Y}}^{\ell}_{jm}(\theta,\phi)$ (see supplementary equations (1-3)). They correspond to the azimuthal ($q=0$), polar ($q=1$), and radial ($q=-1$) components. In the context of Mie scattering, $\bm{\mathcal{Y}}_{\ell m}^{(1)}(\theta,\phi)$ represents electric multipoles, whereas $\bm{\mathcal{Y}}_{\ell m}^{(0)}(\theta,\phi)$ corresponds to magnetic multipoles. In homogeneous media, the radial component is not allowed for electromagnetic fields. Together with Equation (31), we show in the Supplementary Method 1 how to express the azimuthal or polar vector spherical harmonics using the Cartesian multipole expansion of an appropriate current source. We show that a current density given by
@@
where $\mathbf{Y}_{\ell\pm 1,m}^{\ell}(\mathbf{k})=k^{\ell}\bm{\mathcal{Y}}_{\ell\pm 1,m}^{\ell}(\mathbf{k})$. The Cartesian components of the current densities $\mathbf{J}_{\ell m}^{(0,1)}(\mathbf{r})$ correspond to singularities of the scalar wave equation analyzed in the “Spherical and Cartesian multipole expansions in isotropic media” subsection of the “Methods” section. Hence, any component of the vector spherical harmonics can be represented by an appropriate sum of Cartesian multipoles.
II.3.2 From the Cartesian multipole expansion to the vector spherical harmonics
As expected from the findings for the scalar wave equation, the inverse transformation—i.e., writing an arbitrary vector Cartesian multipole expansion as a linear combination of vector spherical harmonics—is not as straightforward, and a complete treatment goes beyond the scope of this research. To understand the issue at hand, one approach is to count the number of vector spherical harmonics at a given order and to compare with the corresponding number of Cartesian multipole moments. As we will see, as in the scalar case, there is an excess of Cartesian multipole moments.
Let us now count the vector spherical harmonics. The azimuthal vector harmonics correspond to a current density given by the vector spherical harmonics $\bm{\mathcal{Y}}_{\ell m}^{\ell}(\theta,\phi)$. In light of Equation (31), there are $12\ell+$ 1 different azimuthal waves of order $\ell>0$ (none for $\ell=0$). By the same argument, there are $12(\ell\pm 1)+$ 1 polar vector harmonics of order $\ell>0$ corresponding to $j=\ell\pm 1$. But the polar harmonics are degenerate, in that both values of $j$ yield the same waves, by virtue of the contraction in Equation (26) (see the Supplementary Method 1.2.3). Thus, in total, there are
@@ II.3.2 From the Cartesian multipole expansion to the vector spherical harmonics
Next, we count the moments for the vector Cartesian multipole expansion. Since there are three possible polarizations ($x$, $y$, and $z$) and $12\ell+$ 1 spherical harmonics of order $\ell$ per polarization, one would expect
@@
different vector spherical harmonics (VSH). The missing harmonics are the degenerate polar harmonics with $j=\ell-1$, and their count is given by 2($12(\ell-1)+$ 1:
$\underbrace{4\ell+4}_{\text{VSH}}+\underbrace{2(\ell-1)+1}_{\text{degenerate}}=\underbrace{6\ell+3}_{\text{Cartesian}}$ (38)
This hints towards a similar issue to the scalar case when going to anisotropic media, as the Laplace operator is responsible for the degeneracy. Another viewpoint is that in anisotropic media, the degeneracy disappears, and new degrees of freedom in wave polarization become available.
To note, we circumvent this issue in the section “Illustration on an inverse problem” by opting for the electric Green’s tensor-valued function instead of vector spherical harmonics. This approach enables a direct connection between the electric current and the vector-valued field.